Please Help me solve...
Ms. Mills invested her $20,000 bonus in two accounts. She took a 4% loss on one investment and made a 12% profit on another investment but ended up breaking even. How much was invested in each account?

Answer by stanbon(75887)

Ms. Mills invested her $20,000 bonus in two accounts. She took a 4% loss on one investment and made a 12% profit on another investment but ended up breaking even. How much was invested in each account?
----------------
Equations:
f + t = 20,000
-0.04f + 0.12t = 0
------------------------
Modify:
Multiply thru the top equation by 4:
Multiply thru the bottom equation by 100::
4f + 4t = 80,000
-4f+12t = 0
------------------------
Add and solve for "t":::
16t = 80,000
t = 5,000 (amt. invested at 12%)
-----
Solve for "f":
f + 5000 = 20,000
f = 15,000 (amt. invested at 4%)
--------------------
Cheers,
Stan H.
================